# Intents, Techniques, and Components: a Unified Analysis of Interaction Authoring Tasks in Data Visualization

### v2

@@ Abstract.
© 2024 All Authors. This v2 supersedes v1 (arXiv:2409.01399v1), which had a different license. Do not use v1.
@@ 1. Introduction
As the support for creating static visual representations of data has matured in visualization grammars and authoring tools (Liu et al., 2018; Ren et al., 2018; Satyanarayan and Heer, 2014; Tsandilas, 2020; Wang et al., 2021), there is a growing interest in enabling easy interaction authoring— adding interactive behavior to static visualizations—through declarative languages and interactive systems (Satyanarayan et al., 2016; Zong et al., 2020; Latif et al., 2022; Sultanum et al., 2021; Mörth et al., 2022). To understand the expressiveness of extant interaction authoring systems and to inform the design of new languages and tools, it is important to have theories and models that describe the task space of interaction authoring in data visualization. While previous research has provided taxonomies and frameworks (Yi et al., 2007; Heer and Shneiderman, 2012; Munzner and Brehmer, 2013; Gotz and Zhou, 2009) to describe tasks in interactive visualization, they primarily focus on how visualizations are used for analytical or communication purposes, rather than how interactivity is composed. In this paper, we address this gap by providing a unified framework that describes interaction authoring tasks at multiple levels of abstraction.
To understand what level of abstraction means in the context of interaction authoring, consider the scenario where an author wants to add interactivity to a static visualization consisting of multiple line graphs showing stock prices of various companies over a period of time. At the lowest level, the author may specify that when the user draws a path, the visualization should update to show only the line marks with trends similar to the path drawn. This concrete specification may be composed using a programming language. At a higher level, this specification can be more abstractly described as “query sketch” (Wattenberg, 2001), which may be provided as a standard template in an authoring tool with graphical user interfaces. At the highest level, the author may have the option to just state that she wants to add filtering support to a recommendation system, which can generate applicable techniques and corresponding low-level implementations for her to choose from.
@@ 2.1. Interaction Authoring
Interaction is a frequently used term in a wide range of literature on HCI and Visualization. These two survey papers (Hornbæk and Oulasvirta, 2017; Dimara and Perin, 2019) introduce seven concepts (i.e., Dialogue, Tool Use, Optimal Behavior, Transmission, Embodiment, Experience, and Control) of interaction. Among the concepts, Dialogue (Hornbæk and Oulasvirta, 2017; Dimara and Perin, 2019) defines interaction as a cycle of communication acts channeled between human (via perception and action) and a system (via input and output). By drawing from Norman’s seven stage model (Sedig and Parsons, 2013), this concept is emphasized with mapping that a user seeks for a way to achieve their intention with tasks that a user translates their goals to a intended system state. In addition, tools or models for constructing interactions (Leiva et al., 2019; Blouin and Jézéquel, 2021; Martin et al., 2022; Beaudouin-Lafon, 2000) introduce how they define interaction and how their interaction building system works. Leiva et al. (Leiva et al., 2019) presents a prototyping tool for touch-based mobile interactions. They define interaction as a set of rules to associate user inputs with system outputs. To elaborate, user inputs (i.e., touches, movements, expressions, voice) captured by devices are translated into system outputs (i.e., UI feedback, underlying operations), resulting in visual outputs through feedback from corresponding operations performed by the underlying system such as databases or servers. To construct such interactions, the tool user employs a first-class object called touch events. the user can manipulate touchable UI elements or program functions to establish interactive behaviors (i.e., touch-stamp, pinch, drag, touch), thereby establishing connections between input and output properties (i.e., input: touch, output: the position change of an item). Similarly, other systems define interactions in the same manner: to support dialogues between a system and its users who engage with a UI to control the underlying system (Blouin and Jézéquel, 2021), to establish causal relationships, such as input events instigating state changes or the propagation of values (Martin et al., 2022), and to describe methods to integrate techniques by delineating the feedback from interactions based on the users’ perspective (Beaudouin-Lafon, 2004). While the existing body of literature and systems depict interaction and the process of constructing it as connecting humans and systems across various stages from input to output, this concept is not limited to data visualization.
@@
The survey paper (Dimara and Perin, 2019) argues the importance of involving Data entity and Data-related intent to the definition of interaction in visualization. For example, the entity for interactive data visualization would be data, computer, and human. It is because visualization is for knowledge construction or sense-making data rather than the easiness or intuition of using a system. In addition, visualization, as a data interface, enables human to manipulate an intangible data sources, which can be linked to data-related user intent. This data-oriented interpretation of interaction is also seen in visualization authoring tools and languages (Bostock et al., 2011; Satyanarayan et al., 2016; Lu et al., 2021; Zong et al., 2020; Satyanarayan et al., 2015).
@@ 2.2. Existing Frameworks and Taxonomies in Interactive Visualization
Describing human behavior at multiple levels of abstraction is a cornerstone of activity theory (Brown et al., 1999; Engeström, 1999), which has been informing research and design in human-computer interaction (Kuutti, 1996; Nardi, 1996, 1998). Gotz and Zhou (Gotz and Zhou, 2009) adopt activity theory to describe four levels of abstraction in visual analytics. Pike et al. (Pike et al., 2009) identify high-level and low-level elements of interaction in terms of both user tasks and visualization features. Roth (Roth, 2013) develop a multi-tier taxonomy for cartographic interaction primitives at multiple levels: goals, objectives, operators, and operands. Brehmer and Munzner (Munzner and Brehmer, 2013) present a multi-level typology that describes the why, how, and what dimensions of user task in visualization. These works emphasize the importance of having a holistic understanding across multiple levels of abstraction, but they all focus on tasks in using a visualization, and does not address the creation of interactive visualizations.
@@
Taxonomies and models of interaction in visualization also exist at individual levels. Yi et al. (Yi et al., 2007) introduce a taxonomy of user intent, by raising a need to connect user goals with interaction techniques. Pike et al. (Pike et al., 2009) also discuss interactions at the high-level (user intents) and the low-level (techniques). In addition, Heer and Shneiderman (Heer and Shneiderman, 2012) present categories of interactive dynamics capturing important user tasks. The works offer an understanding of the design space of interactive visualization at the user intent or techniques level.
@@
The early literature also introduced taxonomies at the individual level. Dix et al. (Dix and Ellis, 1998) present four classes of tasks for incorporating simple interaction into static visualization, Buja et al. (Buja et al., 1996) define three classes of view manipulation along with corresponding tasks for high dimensional data visualization, and Shneiderman (Shneiderman, 2003) suggests seven tasks based on data types. However, these taxonomies only cover different classifications at the technique level, necessitating the consolidation of existing taxonomies. (Schulz et al., 2013; Rubab et al., 2021) At the implementation level, Chi et al. (Chi and Riedl, 1998) offer a taxonomy based on operators and operands. Vega-Lite (Satyanarayan et al., 2016) presents a declarative grammar that describes how to compose interactivity with primitives; however, they focus more on the grammars’ capabilities rather than linking different levels. Animated Vegalite (Zong et al., 2022) discusses unified interaction and animation tasks based on their similarities at high and low levels of abstraction. Reactive Vegalite (Satyanarayan et al., 2015) presents a comprehensive data model demonstrating the entire data transformation process, from user input to output, in interactive data visualization. However, these two works primarily focus on lower levels, minimizing the involvement of human activities such as user intent.
@@ 3. Choice of Abstraction Levels
We decide to perform analysis on interaction authoring tasks at multiple levels of abstraction because current literature has pointed out that a single level of abstraction is insufficient to capture the rich semantics of visualization activities in terms of motivation, context, goals and operations (Pike et al., 2009; Roth, 2013; Munzner and Brehmer, 2013; Gotz and Zhou, 2009). The scenario in the introduction illustrates the importance of supporting interaction authoring at multiple levels of abstraction. Although a few multi-level visualization frameworks exist (e.g., Brehmer and Munzner (Munzner and Brehmer, 2013), Gotz and Zhou (Gotz and Zhou, 2009)), we did not closely follow their level characterizations because of our distinct focus on interaction authoring. In this section, we define the three levels of abstraction (i.e., intent, technique, and component) in our analysis and explain why they are chosen by situating them in the context of existing literature.
@@
Intents. In our framework, “intents“ include “authoring intent“ and “user intent“. Firstly, “authoring intent” is what an author wants to achieve when authoring interaction. We define “authoring intent“ based on “user intent“ from existing literature, and the definition of interaction authoring in the section 2. Yi et al. (Yi et al., 2007) emphasize the importance of linking “user intent“ with techniques, demonstrating the impact of “user intent“ throughout the data exploration process. As a result, we regard their taxonomy as a primary reference when consolidating “user intent“. They elaborate that different interaction designs can be interpreted to achieve the same goal, called “user intent“. For example, consider the following two designs: (1) clicking on a circle in a scatter plot changes its color, and (2) swiping horizontally over a bar chart on a phone fades the bars outside the swiped region into the background. Both interaction designs serve the same intent: to select items of interest. Besides Yi et al.’s taxonomy, Heer and Shneiderman’s interactive dynamics categories (Heer and Shneiderman, 2012) and Brehmer and Munzner’s how level (Munzner and Brehmer, 2013) are also operating at this level of abstraction for “user intent“. The former presents a more extensive list of categories, while the latter not only addresses the user intent but also examines the relationships between user intents and other levels. Taking these interpretations into account, we define “authoring intent” that is to seek a task to translate these “user intents“ to desired system elements. For example, to enable users to select a bar, an author wants to connect a bar and corresponding changes in its visual properties by programming.
@@
Techniques. In our framework, techniques are to achieve a user intent, which is ultermately for the goal of fulfulling authoring intent. Techniques often tightly coupled with the type of items being manipulated. For instance, the two designs in the previous paragraph are examples of two different techniques: point select and range select. The former is about selecting a single item, while the latter is about selecting items within a specific pixel range. For the same reasons associated with the intents, we regard the taxonomies by Yi et al. (Yi et al., 2007) and Heer and Shneiderman (Heer and Shneiderman, 2012) as our primary references. These taxonomies feature representative techniques for each high-level user intent. We consolidate their techniques and add more techniques discovered while analyzing examples.
@@
Components. Finally, components are the building blocks to implement interaction designs. We consider this level similar to the what level in Brehmer and Munzner’s typology (Munzner and Brehmer, 2013). Their account at the what level, however, is not detailed and does not address interaction authoring. In general, there is little work that systematically identifies the primitives in authoring interactive visualizations. A represented relevant work is Vega-Lite (Satyanarayan et al., 2016), which identifies what comprises interactivity from the perspective of declarative specification. Our analysis at this level seeks to provide more details on what inputs and outputs are involved to translate UI events into visual changes.
@@ 4. Analysis Method and Process
The primary goal of our analysis is to identify common categories at each of the three levels and organize these categories across levels. Similar to previous work that focuses on identifying high-level intents (Yi et al., 2007) and low-level primitives (Liu et al., 2018; Thompson et al., 2020) in data visualization, we ground this work in the analysis of real-world interactive visualization examples.
@@
Based on these criteria, we identified 47 real-world interactive visualization examples from 16 online sources, including web applications (e.g., GOV||DNA (Sudox, [n. d.]), ESA Star Mapper (Interactive, [n. d.])), news outlets (e.g., the New York Times (The New York Times, [n. d.]), the Economist (The Economist, [n. d.])), digital publishers (e.g., the Pudding (The Pudding, [n. d.]), Quartz (Quartz, [n. d.])), design galleries of visualization toolkits (e.g., bl.ocks (Popular Blocks, [n. d.])) and systems (e.g., Tableau Public (Tableau, [n. d.]), datasketch (Datasketch, [n. d.])), commercial demos (e.g., heavy.ai (Heavy.ai, [n. d.]), Gramener (Gramener, [n. d.])), and portfolio sites of design agencies (e.g., Column Five Media (Column Five Media, [n. d.])).
@@
Phase 1: We first adopted a top-down approach to apply existing taxonomies to the examples. At the intent level, we tried to assign each unit of analysis a category by consolidating Yi et al.’s taxonomy (Yi et al., 2007), Heer and Shneiderman’s taxonomy (Heer and Shneiderman, 2012), and Brehmer and Munzner’s how classification (Munzner and Brehmer, 2013). At the component level, we analyzed each unit of analysis by labeling the selection primitives in Vega-Lite’s interaction grammar (e.g., type, predicate, domain—range, init, transforms, resolve). At the early stage, All three authors independently coded the examples and met to resolve disagreements. Initially, discussions on a single complex example could take an hour due to differing interpretations of intent categories and selection primitives. However, as we aligned on the concepts, we could review three examples within an hour. After the initial ten examples, two authors continued coding and involved the third only when needed for disagreements.
@@ 5.1. Issues with Directly Using Existing Taxonomies for Interaction Authoring
As mentioned in Section 3, we used three existing taxonomies (Yi et al., 2007; Heer and Shneiderman, 2012; Munzner and Brehmer, 2013) to label intent. While the taxonomies were highly useful in providing conceptual scaffolding in our analysis, we did encounter issues that prevent us from directly applying these existing taxonomies to interaction authoring: focus on user intent instead of authoring intent.
@@
The existing taxonomies describe intent in terms of what users want to achieve. This is evident in (Yi et al., 2007), where the meaning of each intent is explained in the form of “show me …”, and in (Heer and Shneiderman, 2012), where the “categories incorporate the critical tasks that enable iterative visual analysis”. However, to enable users to achieve their goals, authors need to first articulate what they want to achieve in terms of how the target visual objects should change. In other words, what to connect to change the target visual objects. For example, in Figure 2 the user intent here is elaborate using Yi et al.’s taxonomy (Yi et al., 2007), which describes what users want to achieve: to get more information about the station. To support this user intent, authors need to specify how the visibility and position of the tooltip should be updated based on mouse events, as well as how the tooltip content should be updated by retrieve the corresponding data. To do so, they should intend to link the input event, the tooltip, and its visual properties. The authoring intents here are very different from the user intent elaborate.

### v3

@@ Abstract.
© 2025 All Authors. Version 3 (v3) supersedes v1 and v2, which used different license and acknowledgment terms. Please do not use v1 or v2.
@@ 1. Introduction
As the support for creating static visual representations of data has matured in visualization grammars and authoring tools (liu_data_2018; ren_charticulator_2018; satyanarayan_lyra_2014; tsandilas_structgraphics_2020; wang_falx_2021), there is a growing interest in enabling easy interaction authoring— adding interactive behavior to static visualizations—through declarative languages and interactive systems (satyanarayan_vega-lite_2016; zong_lyra_2020; latif_kori_2022; sultanum_leveraging_2021; morth_scrollyvis_2022). To understand the expressiveness of extant interaction authoring systems and to inform the design of new languages and tools, it is important to have theories and models that describe the task space of interaction authoring in data visualization. While previous research has provided taxonomies and frameworks (yi_toward_2007; heer_interactive_2012; munzner_multi-level_2013; gotz_characterizing_2009) to describe tasks in interactive visualization, they primarily focus on how visualizations are used for analytical or communication purposes, rather than how interactivity is composed. In this paper, we address this gap by providing a unified framework that describes interaction authoring tasks at multiple levels of abstraction.
To understand what level of abstraction means in the context of interaction authoring, consider the scenario where an author wants to add interactivity to a static visualization consisting of multiple line graphs showing stock prices of various companies over a period of time. At the lowest level, the author may specify that when the user draws a path, the visualization should update to show only the line marks with trends similar to the path drawn. This concrete specification may be composed using a programming language. At a higher level, this specification can be more abstractly described as “query sketch” (wattenberg_sketching_2001), which may be provided as a standard template in an authoring tool with graphical user interfaces. At the highest level, the author may have the option to just state that she wants to add filtering support to a recommendation system, which can generate applicable techniques and corresponding low-level implementations for her to choose from.
@@ 2.1. Interaction Authoring
Interaction is a frequently used term in a wide range of literature on HCI and Visualization. These two survey papers (hornbaek_what_2017; dimara_what_2019) introduce seven concepts (i.e., Dialogue, Tool Use, Optimal Behavior, Transmission, Embodiment, Experience, and Control) of interaction. Among the concepts, Dialogue (hornbaek_what_2017; dimara_what_2019) defines interaction as a cycle of communication acts channeled between human (via perception and action) and a system (via input and output). By drawing from Norman’s seven stage model (sedig_interaction_2013), this concept is emphasized with mapping that a user seeks for a way to achieve their intention with tasks that a user translates their goals to a intended system state. In addition, tools or models for constructing interactions (leiva_enact_2019; blouin_interacto_2021; martin_causette_2022; beaudouin-lafon_instrumental_2000) introduce how they define interaction and how their interaction building system works. Leiva et al. (leiva_enact_2019) presents a prototyping tool for touch-based mobile interactions. They define interaction as a set of rules to associate user inputs with system outputs. To elaborate, user inputs (i.e., touches, movements, expressions, voice) captured by devices are translated into system outputs (i.e., UI feedback, underlying operations), resulting in visual outputs through feedback from corresponding operations performed by the underlying system such as databases or servers. To construct such interactions, the tool user employs a first-class object called touch events. the user can manipulate touchable UI elements or program functions to establish interactive behaviors (i.e., touch-stamp, pinch, drag, touch), thereby establishing connections between input and output properties (i.e., input: touch, output: the position change of an item). Similarly, other systems define interactions in the same manner: to support dialogues between a system and its users who engage with a UI to control the underlying system (blouin_interacto_2021), to establish causal relationships, such as input events instigating state changes or the propagation of values (martin_causette_2022), and to describe methods to integrate techniques by delineating the feedback from interactions based on the users’ perspective (beaudouin-lafon_designing_2004). While the existing body of literature and systems depict interaction and the process of constructing it as connecting humans and systems across various stages from input to output, this concept is not limited to data visualization.
@@
The survey paper (dimara_what_2019) argues the importance of involving Data entity and Data-related intent to the definition of interaction in visualization. For example, the entity for interactive data visualization would be data, computer, and human. It is because visualization is for knowledge construction or sense-making data rather than the easiness or intuition of using a system. In addition, visualization, as a data interface, enables human to manipulate an intangible data sources, which can be linked to data-related user intent. This data-oriented interpretation of interaction is also seen in visualization authoring tools and languages (bostock_d3_2011; satyanarayan_vega-lite_2016; lu_automatic_2021; zong_lyra_2020; satyanarayan_reactive_2015).
@@ 2.2. Existing Frameworks and Taxonomies in Interactive Visualization
Describing human behavior at multiple levels of abstraction is a cornerstone of activity theory (brown_perspectives_1999; engestrom_activity_1999), which has been informing research and design in human-computer interaction (kuutti_activity_1996; nardi_activity_1996; nardi_context_1998). Gotz and Zhou (gotz_characterizing_2009) adopt activity theory to describe four levels of abstraction in visual analytics. Pike et al. (pike_science_2009) identify high-level and low-level elements of interaction in terms of both user tasks and visualization features. Roth (roth_empirically-derived_2013) develop a multi-tier taxonomy for cartographic interaction primitives at multiple levels: goals, objectives, operators, and operands. Brehmer and Munzner (munzner_multi-level_2013) present a multi-level typology that describes the why, how, and what dimensions of user task in visualization. These works emphasize the importance of having a holistic understanding across multiple levels of abstraction, but they all focus on tasks in using a visualization, and does not address the creation of interactive visualizations.
@@
Taxonomies and models of interaction in visualization also exist at individual levels. Yi et al. (yi_toward_2007) introduce a taxonomy of user intent, by raising a need to connect user goals with interaction techniques. Pike et al. (pike_science_2009) also discuss interactions at the high-level (user intents) and the low-level (techniques). In addition, Heer and Shneiderman (heer_interactive_2012) present categories of interactive dynamics capturing important user tasks. The works offer an understanding of the design space of interactive visualization at the user intent or techniques level.
@@
The early literature also introduced taxonomies at the individual level. Dix et al. (dix_starting_1998) present four classes of tasks for incorporating simple interaction into static visualization, Buja et al. (buja_interactive_1996) define three classes of view manipulation along with corresponding tasks for high dimensional data visualization, and Shneiderman (shneiderman_eyes_2003) suggests seven tasks based on data types. However, these taxonomies only cover different classifications at the technique level, necessitating the consolidation of existing taxonomies. (schulz_design_2013; rubab_examining_2021) At the implementation level, Chi et al. (chi_operator_1998) offer a taxonomy based on operators and operands. Vega-Lite (satyanarayan_vega-lite_2016) presents a declarative grammar that describes how to compose interactivity with primitives; however, they focus more on the grammars’ capabilities rather than linking different levels. Animated Vegalite (zong_animated_2022) discusses unified interaction and animation tasks based on their similarities at high and low levels of abstraction. Reactive Vegalite (satyanarayan_reactive_2015) presents a comprehensive data model demonstrating the entire data transformation process, from user input to output, in interactive data visualization. However, these two works primarily focus on lower levels, minimizing the involvement of human activities such as user intent.
@@ 3. Choice of Abstraction Levels
We decide to perform analysis on interaction authoring tasks at multiple levels of abstraction because current literature has pointed out that a single level of abstraction is insufficient to capture the rich semantics of visualization activities in terms of motivation, context, goals and operations (pike_science_2009; roth_empirically-derived_2013; munzner_multi-level_2013; gotz_characterizing_2009). The scenario in the introduction illustrates the importance of supporting interaction authoring at multiple levels of abstraction. Although a few multi-level visualization frameworks exist (e.g., Brehmer and Munzner (munzner_multi-level_2013), Gotz and Zhou (gotz_characterizing_2009)), we did not closely follow their level characterizations because of our distinct focus on interaction authoring. In this section, we define the three levels of abstraction (i.e., intent, technique, and component) in our analysis and explain why they are chosen by situating them in the context of existing literature.
@@
Intents. In our framework, “intents“ include “authoring intent“ and “user intent“. Firstly, “authoring intent” is what an author wants to achieve when authoring interaction. We define “authoring intent“ based on “user intent“ from existing literature, and the definition of interaction authoring in the section 2. Yi et al. (yi_toward_2007) emphasize the importance of linking “user intent“ with techniques, demonstrating the impact of “user intent“ throughout the data exploration process. As a result, we regard their taxonomy as a primary reference when consolidating “user intent“. They elaborate that different interaction designs can be interpreted to achieve the same goal, called “user intent“. For example, consider the following two designs: (1) clicking on a circle in a scatter plot changes its color, and (2) swiping horizontally over a bar chart on a phone fades the bars outside the swiped region into the background. Both interaction designs serve the same intent: to select items of interest. Besides Yi et al.’s taxonomy, Heer and Shneiderman’s interactive dynamics categories (heer_interactive_2012) and Brehmer and Munzner’s how level (munzner_multi-level_2013) are also operating at this level of abstraction for “user intent“. The former presents a more extensive list of categories, while the latter not only addresses the user intent but also examines the relationships between user intents and other levels. Taking these interpretations into account, we define “authoring intent” that is to seek a task to translate these “user intents“ to desired system elements. For example, to enable users to select a bar, an author wants to connect a bar and corresponding changes in its visual properties by programming.
@@
Techniques. In our framework, techniques are to achieve a user intent, which is ultermately for the goal of fulfulling authoring intent. Techniques often tightly coupled with the type of items being manipulated. For instance, the two designs in the previous paragraph are examples of two different techniques: point select and range select. The former is about selecting a single item, while the latter is about selecting items within a specific pixel range. For the same reasons associated with the intents, we regard the taxonomies by Yi et al. (yi_toward_2007) and Heer and Shneiderman (heer_interactive_2012) as our primary references. These taxonomies feature representative techniques for each high-level user intent. We consolidate their techniques and add more techniques discovered while analyzing examples.
@@
Components. Finally, components are the building blocks to implement interaction designs. We consider this level similar to the what level in Brehmer and Munzner’s typology (munzner_multi-level_2013). Their account at the what level, however, is not detailed and does not address interaction authoring. In general, there is little work that systematically identifies the primitives in authoring interactive visualizations. A represented relevant work is Vega-Lite (satyanarayan_vega-lite_2016), which identifies what comprises interactivity from the perspective of declarative specification. Our analysis at this level seeks to provide more details on what inputs and outputs are involved to translate UI events into visual changes.
@@ 4. Analysis Method and Process
The primary goal of our analysis is to identify common categories at each of the three levels and organize these categories across levels. Similar to previous work that focuses on identifying high-level intents (yi_toward_2007) and low-level primitives (liu_data_2018; thompson_understanding_2020) in data visualization, we ground this work in the analysis of real-world interactive visualization examples.
@@
Based on these criteria, we identified 47 real-world interactive visualization examples from 16 online sources, including web applications (e.g., GOV||DNA (sudox_govdna_nodate), ESA Star Mapper (interactive_esa_nodate)), news outlets (e.g., the New York Times (the_new_york_times_new_nodate), the Economist (the_economist_big_nodate)), digital publishers (e.g., the Pudding (the_pudding_pudding_nodate), Quartz (quartz_quartz_nodate)), design galleries of visualization toolkits (e.g., bl.ocks (popular_blocks_popular_nodate)) and systems (e.g., Tableau Public (tableau_tableau_nodate), datasketch (datasketch_datasketch_nodate)), commercial demos (e.g., heavy.ai (heavyai_1990_nodate), Gramener (gramener_gramener_nodate)), and portfolio sites of design agencies (e.g., Column Five Media (column_five_media_column_nodate)).
@@
Phase 1: We first adopted a top-down approach to apply existing taxonomies to the examples. At the intent level, we tried to assign each unit of analysis a category by consolidating Yi et al.’s taxonomy (yi_toward_2007), Heer and Shneiderman’s taxonomy (heer_interactive_2012), and Brehmer and Munzner’s how classification (munzner_multi-level_2013). At the component level, we analyzed each unit of analysis by labeling the selection primitives in Vega-Lite’s interaction grammar (e.g., type, predicate, domain—range, init, transforms, resolve). At the early stage, All three authors independently coded the examples and met to resolve disagreements. Initially, discussions on a single complex example could take an hour due to differing interpretations of intent categories and selection primitives. However, as we aligned on the concepts, we could review three examples within an hour. After the initial ten examples, two authors continued coding and involved the third only when needed for disagreements.
@@ 5.1. Issues with Directly Using Existing Taxonomies for Interaction Authoring
As mentioned in Section 3, we used three existing taxonomies (yi_toward_2007; heer_interactive_2012; munzner_multi-level_2013) to label intent. While the taxonomies were highly useful in providing conceptual scaffolding in our analysis, we did encounter issues that prevent us from directly applying these existing taxonomies to interaction authoring: focus on user intent instead of authoring intent.
@@
The existing taxonomies describe intent in terms of what users want to achieve. This is evident in (yi_toward_2007), where the meaning of each intent is explained in the form of “show me …”, and in (heer_interactive_2012), where the “categories incorporate the critical tasks that enable iterative visual analysis”. However, to enable users to achieve their goals, authors need to first articulate what they want to achieve in terms of how the target visual objects should change. In other words, what to connect to change the target visual objects. For example, in Figure 2 the user intent here is elaborate using Yi et al.’s taxonomy (yi_toward_2007), which describes what users want to achieve: to get more information about the station. To support this user intent, authors need to specify how the visibility and position of the tooltip should be updated based on mouse events, as well as how the tooltip content should be updated by retrieve the corresponding data. To do so, they should intend to link the input event, the tooltip, and its visual properties. The authoring intents here are very different from the user intent elaborate.